I'm having issues with this value mixture word problem.
The manager of a farmer's market has 350 lb of grain that costs $1.90 per pound. How many pounds of meal costing $0.80 per pound should be mixed with the 350 lb of grain to produce a mixture that costs $1.50 per pound?
So far, I have:
A...........C........V
(350)......(1.9)....(665)
(x)........(1.5)....(1.5x)
____________________________
(350+x)...(.8).....(.8)(350+x)
I can't seem to find a solution and would love a little boost! Thank you so much!

Answer by stanbon(75887)

The manager of a farmer's market has 350 lb of grain that costs $1.90 per pound. How many pounds of meal costing $0.80 per pound should be mixed with the 350 lb of grain to produce a mixture that costs $1.50 per pound?
--------------------------
Equation:
value + value = value
1.9*350 + 0.80x = 1.5(350+x)
-------
Multiply thru by 10 to get:
-------
19*350 + 8x = 15*350+ 15x
----
7x = 4*350
----
x = 4*50
x = 200 lb (amt. of $0.80 per lb meal needed)
====================
Cheers,
Stan H.
======================